A group of space explores from the Marchante Space Corp have landed on an unknown planet. In order to leave they must find the escape velocity needed to get into orbit. The mass of the planet is 0.5 x 1025 kg and has a radius of 6,598 km. Find the needed velocity to get into orbit. Give your answer in: km/s.
Answer #1

Using conservation of the energy (for escape velocity kinetic energy =potential energy at surface of the planet)

mV2 /2=GMm/R

V2=GM/R ..1

here G=6.67X10-11, M=0.5X1025Kg ,R=6.598X106 m

Put these value in equation ..1

V2=6.67X10-11X0.5X1025/6.598X106

V2=0.5055X108

V=sqrt(0.5055X108)

V=0.711X104

V=7.11 km/s answer